# Quantum Otto heat engine and quantum Mpemba effect in quasiperiodic systems

### v1

@@ Abstract
In this paper, a quasiperiodic system with off-diagonal and diagonal modulations is studied. By analyzing the inverse participation ratio and the centroid position of the wave packet, the delocalization-localization phase diagrams in the equilibrium state and the non-equilibrium steady state are obtained respectively. The stability of the system’s initial state on dissipation and the dissipation-driven delocalization-localization transition (or vice versa) are discussed. The dynamic evolution behavior provides evidence for the existence of the quantum Mpemba effect in this dissipation-modulated system. A starting-line hypothesis is proposed, which can be used to explain the occurrence and absence of the quantum Mpemba effect. In addition, the thermodynamic applications of this system have also been studied. The results prove that the localized phase is beneficial to the realization of quantum heater as well. Our work has promoted the understanding of the steady-state phase transitions and dynamic behaviors of dissipation-modulated quasiperiodic systems, and has expanded the existing thermodynamic research on the application of quasiperiodic systems.
@@ I Introduction
Under different physical mechanisms, Anderson localization takes various forms. In quasiperiodic systems, when the quasiperiodic potential potential strength exceeds a critical value, all quantum states become localized Aubry and André (1980); Roati et al. (2008). In one-dimensional quasiperiodic systems with short- (long-) range hoppings Biddle and Sarma (2010); Deng et al. (2019); Roy and Sharma (2021); Biddle et al. (2011); Liu et al. (2021a); Saha et al. (2019); Xu et al. (2021); Tang et al. (2021), and generalized quasiperiodic modulations Sarma et al. (1988); Ganeshan et al. (2015); Yao et al. (2019); Liu and Guo (2018); Wang et al. (2020); Liu et al. (2022); Zhou et al. (2023); Liu et al. (2020a, 2021b); Gonçalves et al. (2022, 2023a, 2023b, 2024a, 2024b); Varma et al. (2016), Anderson localization occurs only as specific energy levels, separated from delocalized levels by mobility edges. This gives the system an intermediate-phase, neither fully delocalized nor fully localized. Recently, the hidden self duality in the system with quasipriodic modulations is discovered Hu et al. (2025a), which further advances the understanding about the Anderson localization and mobility edges Li et al. (2025); Zhou et al. (2025).
Regarding to the quasiperiodic systems, the researches on mobility edges and intermediate phase have enabled applications such as energy current rectification Balachandran et al. (2019); Madhumita and Maiti (2019, 2017) and superradiance light sources Yin et al. (2020). Recently, the attention on the applications of the quasiperiodic systems are being paid to the finite-temperature aspect. It has been found that the quasiperiodic systems with mobility edges interacting with external heat baths can be used to engineer quantum heat engine Chiaracane et al. (2020); Xu and Cheng (2025). When off-diagonal quasiperiodic systems (without mobility edges) are adopted as the working medium in quantum Otto cycles Xu and Cheng (2025), a more diverse set of working modes has been identified-with the quantum heat engine being one such mode, alongside quantum accelerator and quantum heater. These findings motivate us to study when both diagonal and off-diagonal quasi-periodic modulation are introduced simultaneously and this system is used as the working medium of the quantum Otto cycle, what are the similarities and differences between the working mode of this cycle and the result when only off-diagonal modulations are considered, and what is the connection with the localized phase transition at zero temperature?
@@
Besides the thermodynamical applications, in recent years, the burgeoning field of non-Hermitian physics Longhi (2019); Weidemann et al. (2022); Zheng et al. (2024); Liu and Chen (2024); Xu et al. (2022); Liu et al. (2020b); Purkayastha et al. (2021) and the accurate manipulation of quantum coherence in experiments Lin et al. (2022); Liang et al. (2022); Zhang et al. (2023); Zhao et al. (2025) have spurred intense interest in dissipative open quantum systems. Dissipation proves pivotal in inducing both localized and delocalized states, offering a crucial mechanism to decipher electron transport in disordered and homogeneous materials Gopalakrishnan et al. (2017); Rath and Mintert (2020); D. S. Bhakuni and Lev (2023); Longhi (2024a); Liu et al. (2024); Peng et al. (2024a). It can disrupt localization and boost transport Gopalakrishnan et al. (2017); Rath and Mintert (2020); D. S. Bhakuni and Lev (2023), induce mobility edges in systems that previously lacked delocalized-localized transitions Longhi (2024a), and drive quasiperiodic systems into specific delocalized or localized states Liu et al. (2024); Peng et al. (2024b); Hu et al. (2025b), while also modulating the topological properties of topologically trivial and nontrivial insulators Peng et al. (2024a); Feng et al. (2025), and the initial property dependence of the steady state has also been discovered Feng et al. (2025). This motivate us to study the delocalization to localization transformation of the steady state of systems with diagonal and off-diagonal quasi-periodic modulation under dissipation, and discuss the connection between the localization properties of the steady state and the initial properties of the system.
@@
With respect to the dissipation induced delocalization-localization transition, attention are more paid on the the properties of the non-equilibrium steady states. However, the issue that the relaxation rate of systems whose initial state is further from the equilibrium state is actually faster than that of systems whose initial state is closer to the equilibrium state or steady state, namely the quantum Mpempa effect which is a generation of the Mpemba effect at the quantum domain Mpemba and Osborne (1969); Kell (1969); Jeng (2006); Bechhoefer et al. (2021); Teza et al. (2025), remains to be studied. Regarding to the quantum Mpemba effect, it has been theoretically predicted across a diverse range of systems-including quantum dots Chatterjee et al. (2023); Wang and Wang (2024); Graf et al. (2025); Strachan et al. (2024), spin or bosonic systems Longhi (2024b); Nava and Fabrizio (2019); Caceffo et al. (2024); Moroder et al. (2024); Bhore et al. (2025); Westhoff et al. (2025), multi-level systems Carollo et al. (2021); Manikandan (2021); Ivander et al. (2023); Chatterjee et al. (2024), as well as other platforms Chétrite et al. (2021); Holtzman and Raz (2022); Dong et al. (2025); Di Giulio et al. (2025). Experimentally, this effect has also been observed in specific setups: trapped-ion quantum simulators Joshi et al. (2024) and single trapped ions Aharony Shapira et al. (2024); Zhang et al. (2025). Motivated by these outstanding works, we want to know whether the quantum Mpemba effect will occur in this open system with diagonal and off-diagonal quasiperiodic modulations as well. A previous study has shown that if the initial state is localized, the quantum Mpemba effect will occur Dong et al. (2025). Therefore, we want to study whether the quantum Mpemba effect will occur if the initial state is not a localized state as well. Furthermore, we also aim to provide a unified explanation for the appearance and absence of the quantum Mpemba effect.
@@
This work is organized as follows. Section II studies the model with diagonal and off-diagonal modulations and gives out the localization phase diagram. Section III studies the thermodynamical applications of the quantum Otto cycle with the quasiperiodic system being the working medium. Section IV studies the delocalization-localization transition of the steady state after considering the bond dissipations. Section V reveals the phenomenon of quantum Mpemba effect and puts forward a hypothesis to explain the cause of this phenomenon. A summary is presented in Sec. VI.
@@ II Model and localization phase diagram
where $t_{n}=t+t_{1}\cos{(2\pi\alpha n)}$ and $V_{n}=V\cos{(2\pi\alpha n)}$ ($t$ is the unit of energy, $n$ is the site index, and $\alpha=\frac{\sqrt{5}-1}{2}$ is the incommensurate parameter which is the cause of the quasiperiodic modulations.).
@@
The localization phase diagram can be determined from the inverse participation ratio (IPR). Given a normalized wave function $\psi_{j}=\sum^{L}_{n=1}\phi_{j}(n)\hat{c}^{{\dagger}}_{n}|0\rangle$ (Here $j$ is the index of wave function), the corresponding ${\rm IPR}_{j}$ is
@@
It is known that the the $\rm IPR$} of the extended state, critical state, and localized state respectively possess the properties of ${\rm IPR}\approx 0$, $0<{\rm IPR}<1$, and ${\rm IPR}\approx 1$ Biddle and Sarma (2010). Under the given parameters, the extended phase, critical phase and localized phase can be further characterized by the average value of $\rm IPR$} of all wave functions, namely the average inverse participation ratio (MIPR), which is denoted as ${\rm MIPR}=\sum^{L}_{j=1}{\rm IPR}_{j}/L$. Taking system size $L=1597$, the localization phase diagram in $V$-$t_{1}$ parameter space is plotted in Fig. 1, where the color bar denotes $\log_{10}(\rm MIPR)$. In fact, for the $V=0$ case, the earlier studies have proved that the system exits an extended-critical phase transition Liu et al. (2016). We can see from Fig. 1 that under $V=0$, $\log_{10}(\rm MIPR)$ for the extended phase approaching 3-3 and $\log_{10}(\rm MIPR)$ for the critical phase approaches 1-1. Based on this, we can infer that in the case of a finite $V$, the quarter elliptic region with $\log_{10}(\rm MIPR)$ close to 3-3 should be the extended phase. On its outside, $\log_{10}(\rm MIPR)$ is greater than 1-1, so it should be a localized phase.
Compared with the non-Hermitian situation Chen et al. (2022), the recovery of system’s Hermiticity expands the area of the extended phase. To locate the phase boundary of the extended-localized transition, we shall introduce the transformation $r=\sqrt{\frac{V^{2}}{4}+t^{2}_{1}}$ with $\theta={\rm Arctan}(V/t_{1})\in\left[0,2\pi\right]$. Under $L=1597$, the $\log_{10}(\rm MIPR)$ versus $r$ with different $\theta$ are plotted in Fig. 2(a). Intuitively, we can see that for different $\theta$, the $\log_{10}(\rm MIPR)$-$r$ curves characterizing the extended-localized (critical) transition all present a jump at the critical value $r_{c}/t=1$, namely the phase boundary $\sqrt{\frac{V^{2}}{4}+t^{2}_{1}}=t$, which has been plotted as the red dashed line in Fig. 1.
To further verify the above mentioned conclusions we have drawn, we introduce the fraction dimension $D$. Consider a system where the system size $L$ equals the $m$-th Fibnacci number $F_{m}$ and the incommensurate parameter $\alpha$ is replaced by $\alpha=F_{m-1}/F_{m}$. At a specific lattice site $n$, the fraction dimension $D_{n}$ can be derived from the formula
@@
Here, $p_{n}$ represents the probability density. This equation reveals that $D_{n}$ acts as a scaling index. For an extended state, since $p_{n}\sim 1/F_{m}$, we have $D_{n}\sim 1$. In the case of a localized state, $D_{n}\sim 1$ at the sites where the particle occupies, while $D_{n}\rightarrow\infty$ at unoccupied sites. For a critical state, $D_{n}$ lies within a finite interval $\left[D_{min},D_{max}\right]$. Thus, the minimum value $D_{min}$ directly reflects the nature of a given wave function $\psi_{j}$. Specifically, $D_{min}\rightarrow 0$ indicates a localized state, $0<D_{min}<$ points to a critical state, and $D_{min}\rightarrow 1$ signifies an extended state. For generality, we use the average of $D_{min}$ over all states, defined as $\overline{D_{min}}=\sum^{L}_{j=1}D^{j}_{min}/L$. In the limit of $1/m\rightarrow 0$ (extrapolation limit), this average helps distinguish different phases. We select representative parameter points in various phases to compute $\overline{D_{min}}$. As depicted in Fig. 2(b), we observe that the relevant $D_{min}$ approaches $1$ at the parameter sites $\left(r,\theta\right)=\left(0.5t,\frac{\pi}{4}\right)$ and $\left(r,\theta\right)=\left(0.8t,\frac{\pi}{3}\right)$. This observations confirm that the system resides in the extended phase under these conditions. As anticipated, the corresponding $D_{min}$ values fall within the interval $\left(0,1\right)$ in the thermodynamic limit when $\left(r,\theta\right)=\left(1.5t,0\right)$ and $\left(r,\theta\right)=\left(2t,0\right)$. These results clearly reveal the systems’s critical-phase characteristics in such parameter regimes. When $\left(r,\theta\right)=\left(1.5t,\frac{\pi}{3}\right)$ and $\left(r,\theta\right)=\left(2t,\frac{\pi}{6}\right)$, the associated $D_{min}$ extrapolates to 0. This outcome identifies that the system is in the localized phase at these parameter points.
III Thermodynamical applications
In recent years, some researches have been carried out on the thermodynamical applications of quasi-periodic systems Chiaracane et al. (2020); Xu and Cheng (2025), leading to the identification of their applications in aspects such as quantum heat engine, quantum heater, and quantum accelerator. Specifically, it has been demonstrated that the extended phase plays a role in sustaining the operational mode of heat engine, whereas the critical phase is conducive to maintaining the working modes of quantum heater and quantum accelerator. Within this section, we focus on investigating the thermodynamic applications of this generalized Aubry-André (AA) model with both diagonal and off-diagonal quasiperiodic modulations. Our primary objectives include verifying whether the extended phase of this model still contributes to preserving the working mode of quantum heat engines, as well as determining which specific working mode the localized phase favors.
Figure 3: (Color Online) Schematic illustration of the four-stroke quantum heat cycle process. $T_{h}$ and $T_{c}$ stand for the high-temperature and low-temperature heat sources, respectively. The working medium employed is the generalized AA model with diagonal and off-diagonal quasiperiodic modulations. $r^{i}$ and $r^{f}$ represent the systematic parameters of the corresponding Hamiltonians. The system size is $L=610$.
We leverage this quasiperiodic system as the working medium to construct a quantum heat cycle process. The schematic of this cycle is shown in Fig. 3. This quantum heat cycle can be regarded as the quantum counterpart Quan et al. (2007); Niskanen et al. (2007); Abah et al. (2012); Deng et al. (2013); del Campo et al. (2014); Roßnagel et al. (2016); Zheng et al. (2016); Kosloff and Rezek (2017); Deffner and Campbell (2019); Campisi and Fazio (2016); Buffoni et al. (2019); Solfanelli et al. (2020); Zhang et al. (2022) of the classical Otto cycle. The first (④ $\rightarrow$ ①) and third (② $\rightarrow$ ③) strokes occur under thermal contact with high-temperature ($T_{c}$) heat reservoirs. These strokes proceed without external driving or particle exchange. The second (① $\rightarrow$ ②) and fourth (③ $\rightarrow$ ④) strokes are adiabatic from a thermodynamic viewpoint (meaning thermal isolation), yet they might not be strictly adiabatic in the quantum-mechanical sense, since quantum transitions can occur during the evolution [82].
In the first stroke, the working medium, characterized by the Hamiltonian $H(r^{i})$, will eventually reach thermal equilibrium and relax into a Gibbs state. The density matrix $\rho_{1}$ of this state is given by $\rho_{1}=\frac{e^{-\beta_{h}H(r^{i})}}{Z_{1}}$, where $\beta_{h}=\frac{1}{k_{b}T_{h}}$ ($k_{b}$ is Boltzmann constant) and $Z_{1}={\rm Tr}\left[e^{-\beta_{h}H(r^{i})}\right]$ is the partition function. Thus, at thermal equilibrium, the system energy is $E_{1}={\rm Tr}\left[\rho_{1}H(r^{i})\right]$. During the second stroke, the Hamiltonian parameter is switched from $r^{i}$ to $r_{f}$. In this stroke, only work is done, and no heat exchange takes place. The Gibbs state $\rho_{2}$ of this stroke remains unchanged, i.e, $\rho_{2}=\rho_{1}$, but the energy $E_{2}$ becomes $E_{2}={\rm Tr}\left[\rho_{2}H(r)\right]$. In the third stroke, the medium with $r=r^{f}$ comes into contact with a heat source having $\beta_{c}=\frac{1}{k_{b}T_{c}}$. As a result, the Gibbs state $\rho_{3}=\frac{e^{-\beta_{c}H(r^{f})}}{Z_{2}}$, with the partition function $Z_{2}={\rm Tr}\left[e^{-\beta_{c}H(r^{f})}\right]$. Then, the energy of the medium is $E_{3}={\rm Tr}\left[\rho_{3}H(r^{f})\right]$. The fourth stroke can be regarded as the thermal annealing process. In this process, the hopping parameter in the medium is changed back from $r^{f}$ to $r^{i}$. However, the Gibbs state $\rho_{4}$ stays as $\rho_{4}=\rho_{3}$. Hence, the energy of the medium becomes $E_{4}={\rm Tr}\left[\rho_{4}H(r^{i})\right]$.
Figure 4: (Color Online) Working modes of the four-stroke cycle. (a) $r^{f}=0.5t$ and $\theta=\pi/6$. (b) $r^{f}=0.5t$ and $\theta=\pi/4$. (c) $r^{f}=0.9t$ and $\theta=\pi/6$. (d) $r^{f}=0.9t$ and $\theta=\pi/4$. (e) $r^{f}=1.1t$ and $\theta=\pi/6$. (f) $r^{f}=1.1t$ and $\theta=\pi/4$. The blue regions represent the Heat engine. The yellow regions denote the Accelerator and the brown regions denote the Heater. Other parameters are $L=610$ and $k_{b}T_{c}=0.1t$.
After the working medium undergoes this complete cycle, we can determine the heat absorbed $Q_{h}$ ($Q_{h}=E_{1}-E_{4}$) from the $T_{h}$ source, the heat released $Q_{c}$ ($Q_{c}=E_{3}-E_{2}$) to the $T_{c}$ source, and the net work $W$ done by the working medium, where $W=Q_{h}+Q_{c}$. Importantly, the heat cycle process abides by the Clausius inequality, a cornerstone of thermodynamics. Depending on the values of $Q_{h}$, $Q_{c}$, and $W$, the engine using the extended-localized (critical) quasiperiodic system as the working medium exhibits different modes Buffoni et al. (2019); Solfanelli et al. (2020):
  1. 1.
    Heat engine: $Q_{h}>0$, $Q_{c}<0$, and $W>0$;
  2. 2.
    Refrigerator: $Q_{h}<0$, $Q_{c}>0$, and $W<0$;
  3. 3.
    Heater: $Q_{h}<0$, $Q_{c}<0$, and $W<0$;
  4. 4.
    Accelerator: $Q_{h}>0$, $Q_{c}<0$, and $W<0$.
For a system with size $L=610$ and $k_{b}T_{c}=0.1t$, after analyzing the values of $Q_{h}$, $Q_{c}$ and $W$, the corresponding working modes of the four-stroke cycle under different $r^{f}$ and $\theta$ are plotted in Figs. 4(a)-4(f), respectively. It can be observed that this cycle exhibits a diverse range of working modes. The blue regions represent the Heat engine mode. The yellow regions correspond to the Accelerator mode, while the brown regions indicate the Heater working mode. It is noted from Figs. 4(a)-4(d) that when $r^{f}$ and $r^{i}$ are below the critical value $r_{c}$, the Heat engine mode appears. This suggests that similar to the results of the critical-extended quasi-periodic system Xu and Cheng (2025), for the quasi-periodic system under a finite $V$, its extended phase is favorable for the design of a quantum heat engine as well. As $r^{f}$ increases and surpasses the critical values $r_{c}=t$, the system entries the localized phase. From Figs. 4(e) and 4(f), we can see that regions representing the Heater mode expand. This indicates that the localized phase is more conducive to the realization of a quantum heater. Additionally, it is clear that there are extensive parameter ranges corresponding to the Accelerator mode, regardless of whether the system is in the extended phase or the localized phase. Combining with previous research on extended-critical quasiperiodic system Xu and Cheng (2025), we draw a conclusion that all the extended, critical and localized phases facilitate the realization of a quantum accelerator, and both the localized and critical phases are conducive to the realization of quantum heater.
Moreover, the results shown in Fig. 4 offer strategies for regulating the working mode of the four-stroke cycle. The transition between different working modes can be achieved by tuning $r^{f}$, $r^{i}$, and $k_{b}T_{h}$. For example, when $r^{f}$ is much smaller than $r_{c}$ ( refer to Figs. 4(a) and 4(b)), or vice versa, there are two working modes, i.e., Accelerator and Heat engine. By tuning $r^{i}$ and $k_{b}T_{h}$, the Otto cycle can switch between the two working modes. When $r^{f}$ is close to but still less than $r^{c}$ and $k_{b}T_{h}$ is small (refer to Figs. 4(c) and 4(d)), there are three working modes. Thus, the cycle’s working mode will change from Heater to Accelerator and then to Heat engine when we gradually increase $r^{i}$. When $r^{f}$ is larger than $r_{c}$, the cycle can be toggled between Heater and Accelerator modes by tuning $r^{i}$ or $k_{b}T_{h}$.
IV Dissipation driven delocalization-localization transition
In the preceding section, we conducted a systematic investigation into the thermodynamic characteristics of this quasi-periodic system under thermal equilibrium conditions. Our findings revealed two key insights: firstly, the system’s thermodynamic behaviors are tied to its initial properties; secondly, the system holds potential application value for developing quantum heat engine, heater, and accelerator. We also analyzed approaches to regulate these three working modes. In addition to the thermodynamic aspect, recent research efforts have yielded conflicting perspectives regarding the steady-state delocalization-localization transition under dissipative modulation Longhi (2024a); Liu et al. (2024); Peng et al. (2024b); Hu et al. (2025b). By employing the dissipation, one can precisely control the delocalization and localization properties of the system under non-equilibrium steady states. Therefore, it motivates us to study the delocalization and localization properties of the steady state after introducing the dissipation.

### v3

# Quantum Mpemba effect in quasiperiodic systems

@@ Abstract
We study a one-dimensional quasiperiodic tight-binding model with simultaneous off-diagonal (hopping) and diagonal (onsite) modulations. Using the inverse participation ratio and the wave-packet centroid, we construct localization–delocalization phase diagrams for both equilibrium and nonequilibrium steady states. We analyze the robustness of initial-state properties under dissipation and characterize dissipation-induced localization–delocalization transitions (and their reversals) in detail. Trace-distance dynamics provide evidence for a quantum Mpemba effect: states prepared farther from the steady state can relax faster than states initialized closer to it. We propose a starting-line hypothesis that explains the presence or absence of this effect across parameter regimes. These results advance the understanding of steady-state phase transitions and relaxation dynamics in dissipatively driven quasiperiodic systems.
@@ I Introduction
Under different physical mechanisms, Anderson localization takes various forms. In quasiperiodic systems, when the quasiperiodic potential strength exceeds a critical value, all quantum states become localized Aubry and André (1980); Roati et al. (2008). In one-dimensional quasiperiodic systems with short- (long-) range hoppings Biddle and Sarma (2010); Deng et al. (2019); Roy and Sharma (2021); Biddle et al. (2011); Liu et al. (2021); Saha et al. (2019); Xu et al. (2021); Tang et al. (2021) and generalized quasiperiodic modulations Sarma et al. (1988); Ganeshan et al. (2015); Wang et al. (2020); Zhou et al. (2023); Liu et al. (2020); Gonçalves et al. (2023, 2024); Varma et al. (2016) , Anderson localization occurs only as specific energy levels, separated from delocalized level by mobility edges. This gives the system an intermediate-phase, neither fully delocalized nor fully localized. Recently, the hidden self duality in the system with quasiperiodic modulations is discovered Hu et al. (2025a), which further advances the understanding about the Anderson localization and mobility edges Li et al. (2026); Zhou et al. (2025).
@@
As a powerful framework for addressing dissipation Breuer and Petruccione (2007); Ashida et al. (2020), the burgeoning field of accurate manipulation of quantum coherence in experiments Lin et al. (2022); Liang et al. (2022); Zhang et al. (2023); Zhao et al. (2025) have spurred intense interest in dissipative open quantum systems. Dissipation proves pivotal in inducing both localized and delocalized states, offering a crucial mechanism to decipher electron transport in disordered and homogeneous materials Gopalakrishnan et al. (2017); Rath and Mintert (2020); Bhakuni et al. (2024); Longhi (2024); Liu et al. (2024) . It can disrupt localization and boost transport Gopalakrishnan et al. (2017); Rath and Mintert (2020); Bhakuni et al. (2024), induce mobility edges in systems that previously lacked delocalized-localized transitions Longhi (2024), and drive quasiperiodic systems into specific delocalized or localized states Liu et al. (2024); Hu et al. (2025b), while also modulating the topological properties of topologically trivial and nontrivial insulators Feng et al. (2025) , and the initial property dependence of the steady state has also been discovered Feng et al. (2025). This motivate us to study the delocalization to localization transformation of the steady state of systems with diagonal and off-diagonal quasi-periodic modulation under dissipation, and discuss the connection between the localization properties of the steady state and the initial properties of the system.
@@
In studies of dissipation-induced delocalization–localization transitions, prior work has largely concentrated on the properties of nonequilibrium steady states. Indeed, the possibility that states initialized farther from the steady (equilibrium) state may relax faster than those prepared closer to it—an instance of the quantum Mpemba effect, the quantum counterpart of the classical phenomenon—has also attracted considerable attention Mpemba and Osborne (1969); Kell (1969); Jeng (2006); Bechhoefer et al. (2021); Teza et al. (2026). Regarding to the quantum Mpemba effect, it has been theoretically predicted across a diverse range of systems, including quantum dots Chatterjee et al. (2023); Wang and Wang (2024); Graf et al. (2025); Strachan et al. (2025), spin or bosonic systems Nava and Fabrizio (2019); Moroder et al. (2024); Bhore et al. (2025); Westhoff et al. (2025), multi-level systems Carollo et al. (2021); Manikandan (2021); Ivander et al. (2023); Chatterjee et al. (2024), as well as other platforms Holtzman and Raz (2022); Dong et al. (2025); Wei et al. (2025); Xu et al. (2025); Turkeshi et al. (2025). A theoretical framework for the origin of the Mpemba effect in closed many-body quantum systems, together with concise reviews and perspectives on its quantum variants, has been presented Rylands et al. (2024); Ares et al. (2025). Experimental observations have been reported in specific platforms, including trapped-ion quantum simulators Ares et al. (2023), consistent with theoretical predictions, and single trapped-ion systems Aharony Shapira et al. (2024); Zhang et al. (2025). Motivated by recent advances, we examine whether the quantum Mpemba effect can occur in an open, quasiperiodic system with simultaneous diagonal and off-diagonal modulations. While earlier work established the effect for localized initial conditions Dong et al. (2025), here we consider nonlocalized preparations and delineate the conditions under which the effect, if any, survives. We further develop a unified framework that explains both the occurrence and the breakdown of the quantum Mpemba effect.
@@
This work is organized as follows. Section II studies the model with diagonal and off-diagonal modulations and gives out the localization phase diagram. Section III studies the delocalization-localization transition of the steady state after considering the bond dissipations. Section IV reveals the phenomenon of quantum Mpemba effect and puts forward a hypothesis to explain the cause of this phenomenon. A summary is presented in Sec. V.
@@ II Model and localization phase diagram
[…]
@@
The localization phase diagram can be determined from the inverse participation ratio (IPR). Given a normalized wave function $|\psi_{j}\rangle=\sum^{L}_{n=1}\phi_{j}(n)\hat{c}^{{\dagger}}_{n}|0\rangle$ (Here $j$ is the index of wave function), the corresponding ${\rm IPR}_{j}$ is
@@
It is known that the $\rm IPR$} of the extended, critical, and localized state, respectively, possess the properties of ${\rm IPR}\approx 0$, $0<{\rm IPR}<1$, and ${\rm IPR}\approx 1$ Biddle and Sarma (2010). Under the given parameters, the extended, critical, and localized phase can be further characterized by the average value of $\rm IPR$} of all wave functions, namely the mean inverse participation ratio (MIPR), denoted as ${\rm MIPR}=\sum^{L}_{j=1}{\rm IPR}_{j}/L$. With system size $L=1597$, the localization phase diagram in $V$-$t_{1}$ parameter space is plotted in Fig. 1, where the color bar denotes $\log_{10}(\rm MIPR)$. In fact, for the $V=0$ case, the earlier studies have proved that the system exhibits an extended-critical phase transition Liu et al. (2016). We can see from Fig. 1(a) that under $V=0$, $\log_{10}(\rm MIPR)$ for the extended phase approaching 3-3 and $\log_{10}(\rm MIPR)$ for the critical phase approaches 1.6-1.6. Accordingly, for finite $V$, the quarter-elliptic region in parameter space where $\log_{10}(\rm MIPR)\approx-3$ is identified with the extended (delocalized) phase, whereas outside this region $\log_{10}(\rm MIPR)>-1$ indicates localization.
To locate the phase boundary of the extended-localized transition, we shall introduce the transformation $r=\sqrt{V^{2}/4+t^{2}_{1}}$ and $\theta={\rm Arctan}(V/t_{1})$ ($\in\left[0,2\pi\right]$). Under $L=1597$, the $\log_{10}(\rm MIPR)$ versus $r$ with different $\theta$ are plotted in Fig. 2(a). Intuitively, we can see that for different $\theta$, the $\log_{10}(\rm MIPR)$-$r$ curves characterizing the extended-localized (critical) transition all present a jump at the critical value $r_{c}/t=1$, namely the phase boundary $\sqrt{V^{2}/4+t^{2}_{1}}=t$ extracted from the numerical results, which has been plotted as the red dashed line in Fig. 1.
To further verify the above mentioned conclusions, we introduce the fraction dimension $D$. Consider a system where the system size $L$ equals the $m$-th Fibonacci number $F_{m}$ and the incommensurate parameter $\alpha$ is replaced by $\alpha=F_{m-1}/F_{m}$. At a specific lattice site $n$, the fraction dimension $D_{n}$ can be derived from the formula
@@
Here, $p_{n}$ represents the probability density. This equation reveals that $D_{n}$ acts as a scaling index. For an extended state, since $p_{n}\sim 1/F_{m}$, we have $D_{n}\sim 1$. In the case of a localized state, $D_{n}\sim 1$ at the sites where the particle occupies, while $D_{n}\rightarrow\infty$ at unoccupied sites. For a critical state, $D_{n}$ lies within a finite interval $\left[D_{min},D_{max}\right]$. Thus, the minimum value $D_{min}$ directly reflects the nature of a given wave function $|\psi_{j}\rangle$. Specifically, $D_{min}\rightarrow 0$ indicates a localized state, $0<D_{min}<1$ points to a critical state, and $D_{min}\rightarrow 1$ signifies an extended state. For generality, we use the average of $D_{min}$ over all states, defined as $\overline{D_{min}}=\sum^{L}_{j=1}D^{j}_{min}/L$. In the limit of $1/m\rightarrow 0$ (extrapolation limit), this average helps distinguish different phases. We select representative parameter points in various phases to compute $\overline{D_{min}}$. As depicted in Fig. 2(b), we observe that the relevant $D_{min}$ approaches $1$ at the parameter sites $\left(r,\theta\right)=\left(0.5t,\frac{\pi}{4}\right)$ and $\left(r,\theta\right)=\left(0.8t,\frac{\pi}{3}\right)$. These observations confirm that the system resides in the extended phase under these conditions. As anticipated, the corresponding $D_{min}$ values fall within the interval $\left(0,1\right)$ in the thermodynamic limit when $\left(r,\theta\right)=\left(1.5t,0\right)$ and $\left(r,\theta\right)=\left(2t,0\right)$. These results clearly reveal the critical-phase characteristics of the system in such parameter regimes. When $\left(r,\theta\right)=\left(1.5t,\frac{\pi}{3}\right)$ and $\left(r,\theta\right)=\left(2t,\frac{\pi}{6}\right)$, the associated $D_{min}$ extrapolates to 0. This outcome identifies that the system is in the localized phase at these parameter points.
III Dissipation driven delocalization-localization transition